OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter?

OpenStudy (anonymous):

R = 4d - 3 R = 69 69 = 4d-3 4d = 66 d = 16.5 years

OpenStudy (anonymous):

oh ok thankz

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Let daughter's age = x Reggie age be: y y = 4x - 3 69 = 4x - 3 4x = 72 x = 18 Daughter is 18 years old..

OpenStudy (anonymous):

@ramya25 check your solution once again..

OpenStudy (anonymous):

@waterineyes thankz

OpenStudy (anonymous):

It is ok dear..

OpenStudy (anonymous):

thanks @waterineyes

OpenStudy (anonymous):

@ramya25 did you find your mistake??

OpenStudy (anonymous):

yes thanks for correcting lol

OpenStudy (anonymous):

If you know it then it is very nice.. Take care of signs next time.. See, we always know how to solve the problem but while solving we commit silly mistakes of plus minus signs that make our whole solution wrong... Be careful while solving...